### v3

# Augmenting Lateral Thinking in Language Models with Humor and Riddle Data for the BRAINTEASER Task

Abstract
The SemEval 2024 BRAINTEASER task challenges language models to perform lateral thinking—a form of creative, non-linear reasoning that remains underexplored in NLP. The task comprises two subtasks, Sentence Puzzle and Word Puzzle, requiring models to defy conventional commonsense associations.
We present a system that fine-tunes DeBERTaV3 using HuggingFace’s AutoModelForMultipleChoice architecture. We augment the provided training data with two additional sources: (1) a humor-style question-answering dataset generated via GPT-4 prompting, and (2) the RiddleSense dataset. This data augmentation strategy is motivated by the observation that humor and riddles share the lateral reasoning structure required by the task. Our best system achieves 92.5% overall accuracy on the Sentence Puzzle subtask and 80.2% on the Word Puzzle subtask, ranking 6th out of 31 teams and 10th out of 23 teams, respectively. We further show that the choice of task formulation matters: framing the problem as multiple-choice rather than sequence classification yields a 10-point accuracy improvement with the same base model. Our analysis reveals that data augmentation with humor and riddle data is particularly effective for sentence-level lateral reasoning, while word-level puzzles remain a harder challenge.
Augmenting Lateral Thinking in Language Models with Humor and Riddle Data for the BRAINTEASER Task
Mina Ghashamithanks: Both authors contributed equally to this work. Amazon Web Services ghashami@amazon.com          Soumya Smruti Mishra11footnotemark: 1 Amazon Web Services soumish@amazon.com
1 Introduction
The success of language models has inspired the Natural Language Processing community to attend to tasks that require implicit and complex reasoning. Human reasoning encompasses two broad categories: lateral and vertical thinking. Lateral thinking demands unconventional, creative reasoning that deviates from traditional logical processes and has received relatively little attention from the NLP community. Vertical thinking, on the other hand, relies on logical, step-by-step reasoning and has been the focus of considerable research in recent years.
Jiang et al. (2023) introduced the BRAINTEASER dataset as a benchmark for evaluating lateral thinking in question-answering systems. The SemEval 2024 BRAINTEASER task (Jiang et al., 2024) derives a focused challenge from this dataset, comprising two subtasks: Sentence Puzzle, which requires reinterpreting sentence-level meaning, and Word Puzzle, which requires reasoning about orthographic or phonetic properties of words. In both cases, models must go beyond conventional commonsense associations and recognize non-obvious interpretations—a capability that most existing NLP benchmarks, which focus on vertical (logical, step-by-step) reasoning, do not test.
Our system uses pre-trained models—BERT (Devlin et al., 2019) and DeBERTaV3 (He et al., 2023)—through HuggingFace’s (Wolf et al., 2020) AutoModelForMultipleChoice and AutoModelForSequenceClassification architectures. We mix Sentence and Word Puzzle training datasets to broaden exposure to different types of lateral thinking challenges. Additionally, we augment the training data with a humor/jokes dataset generated by GPT-4 (OpenAI et al., 2024) and the RiddleSense (Lin et al., 2021) dataset, both of which require creative, non-literal reasoning similar to the target task.
Our experiments reveal several findings. First, the AutoModelForMultipleChoice formulation substantially outperforms AutoModelForSequenceClassification, confirming that task-appropriate architecture design matters even with strong pre-trained representations. Second, augmenting training data with humor and riddle datasets consistently improves performance on the Sentence Puzzle. Third, the same augmentation strategy does not transfer to the Word Puzzle, suggesting that sentence-level and word-level lateral reasoning require different data characteristics. Our system ranked 6th out of 31 teams on the Sentence Puzzle and 10th out of 23 teams on the Word Puzzle.
Our main contributions are as follows:
  1. 1.
    We show that the multiple-choice architecture (AutoModelForMultipleChoice) significantly outperforms the sequence classification formulation for this task, yielding a 10-point accuracy gain with the same BERT base model.
  2. 2.
    We demonstrate that augmenting training data with GPT-4-generated humor QA data and the RiddleSense dataset improves lateral thinking performance, achieving 92.5% accuracy on the Sentence Puzzle.
  3. 3.
    We provide an analysis of failure modes and the differential effect of data augmentation on sentence-level versus word-level lateral reasoning.
Our code and data will be available at https://github.com/soumyasmruti/semeval-2024-brainteaser after cleaning and de-anonymization.
2 Background
The task involves two types of brain teasers: Sentence Puzzle and Word Puzzle. In Sentence Puzzle, the input is a sentence-based question that defies commonsense, with multiple-choice answers. For instance, "A man shaves everyday, yet keeps his beard long." The choices include "He is a barber," "He wants to maintain his appearance," and so on. The Word Puzzle involves a word-based teaser, like "What part of London is in France?" with choices focusing on letters in the words (e.g., "The letter N"). The output in both cases is the selection of the correct choice that represents lateral thinking.
In order to counter the potential for Large Language Models (LLMs) memorizing solutions, BRAINTEASER (Jiang et al., 2023) incorporates two novel methods of puzzle generation: semantic and context reconstruction. These techniques generate variations of puzzles that preserve the core challenge of overturning conventional commonsense reasoning without altering the fundamental nature of the puzzles. This approach is aimed at enhancing the robustness of the puzzles against the memorization capabilities of LLMs, ensuring that the puzzles continue to effectively test the models’ ability to engage in lateral thinking by challenging ingrained commonsense assumptions. This is to ensure the model is evaluating reasoning ability rather than memorization.
Systems are evaluated based on two accuracy metrics: Instance-based Accuracy, considering each question (original and adversarial) as a separate instance, and Group-based Accuracy, where a system must correctly solve all questions in a group (original and its adversarial versions) to score.
3 Related Work
We can broadly categorize the reasoning landscape of language models into two groups. The first, is ‘commonsense reasoning‘, also known as ‘vertical reasoning‘. This refers to the ability to make deductions based on everyday knowledge. The second category is ‘lateral reasoning‘; i.e. a creative problem-solving approach that involves looking at situations from unconventional perspectives.
Researchers have explored various approaches to endow LLMs with commonsense reasoning abilities (Rae et al., 2021). One prominent approach is the use of knowledge graphs, which represent structured knowledge in the form of entities and their relationships (Ilievski et al., 2021). Wang et al. (2021) proposed a method for incorporating commonsense knowledge from ConceptNet (Speer et al., 2018) into language models, leading to improved performance on commonsense reasoning tasks.
Another approach involves fine-tuning pre-trained LLMs on commonsense reasoning datasets. Huang et al. (2019) introduced the COSMOS QA dataset, which consists of multiple-choice questions that require commonsense reasoning. They showed that fine-tuning pre-trained LLMs on this dataset can significantly improve their commonsense reasoning capabilities.
Researchers have also investigated the use of prompting techniques to elicit commonsense reasoning from LLMs without explicit fine-tuning. Zhou et al. (2022) proposed a method called “Conditional Prompt-Tuning” that enables LLMs to perform commonsense reasoning by conditioning on carefully designed prompts. In another work, Wei et al. (2022) showed that chain-of-thought prompting can unlock LLM reasoning ability via effective prompting techniques.
There has not been extensive research on lateral thinking in LLMs. Recently, Xie et al. (2023) proposed OlaGPT, a cognitive architecture framework that summarizes various methods of human reasoning into Chain-of-Thought (CoT) templates to improve LLM reasoning.
Data Augmentation for Reasoning.
Data augmentation has been widely used to improve model robustness in NLP tasks. In the context of reasoning, augmenting training data with related but distinct datasets has shown promise for improving generalization. Our work extends this idea by using humor and riddle datasets as augmentation sources, motivated by the structural similarity between jokes, riddles, and lateral thinking puzzles—all require reinterpreting familiar concepts in unexpected ways.
Overall, while LLMs have shown flashes of non-linear, exploratory thinking on some benchmarks, lateral thinking as a cognitive process remains an open challenge.
4 Methodology
In this section, we describe different methods and approaches we employed in solving the BrainTeaser puzzle.
4.1 Sequence Classification with BERT
In this approach, we enhanced the performance of a sequence classification model through the instruction fine-tuning process. We leveraged the powerful contextual embeddings provided by BERT (Devlin et al., 2019). Our methodology involved initializing the model with pre-trained BERT weights and employing the streamlined ‘AutoModelForSequenceClassification‘ class from the Hugging Face Transformers library, which linearly projects the embedding from the language model encoder to each document into the class logits for that document. We instructed the model with selecting the most appropriate answer from a set of four choices provided alongside a given question. Despite the meticulous fine-tuning process our experimental results revealed sub-optimal performance.
4.2 MultipleChoice QA with BERT and DeBERTa
We leveraged the versatile ‘AutoModelForMultipleChoice‘ architecture from Hugging Face’s library, which integrates a pre-trained transformer model with a specialized classification head. This architecture was pivotal in adapting the model for our multiple-choice task, which involved combining both Word Puzzle and Sentence Puzzle datasets to diversify our training data.
To ensure optimal performance, we split our training data into separate training and validation sets. Throughout the training process, we utilized the validation set to fine-tune hyperparameters, ensuring the model’s efficacy.
The AutoModelForMultipleChoice architecture comprises a pre-trained base transformer augmented with a classification head. This head, typically consisting of neural network components such as linear layers and activation functions, enables the model to make informed multiple-choice predictions.
Our model initialization involved embedding pre-trained DeBERTa representations, followed by further training on the designated training dataset. This approach facilitated the model’s adaptation to our specific task requirements, ultimately enhancing its performance.
4.2.1 Augmenting with RiddleSense and Humor Data
Next, we decided to use two additional data sources to augment our training data. This was with the aim of expanding the diversity of our dataset, enriching it with a wide range of humor styles, scenarios, and perspectives. This augmentation not only increases the robustness and variety of our model but also enhances its adaptability to different contexts. We utilized the public Riddlesense dataset as well as creating humor style data by prompting GPT 4.
The RiddleSense dataset consists of riddles—questions presented in a cryptic or metaphorical manner that challenge the reader to find a clever or unexpected answer. Riddles share a key structural property with brain teasers: they require the solver to abandon default commonsense interpretations in favor of creative alternatives.
To create the humor-style QA data, we prompted GPT-4 to generate joke-format multiple-choice questions. Jokes frequently rely on lateral thinking, as punchlines subvert the listener’s expectations by reinterpreting the setup in an unexpected way. This structural parallel to the BRAINTEASER task motivates our use of humor data as an augmentation source. The details of the data generation process are provided in Appendix A.
We then used the same AutoModelForMultipleChoice architecture and trained the model on augmented training data.
5 Experimental setup
5.1 Datasets Description
Sentence Puzzle Word Puzzle
Dataset Train Validation Test Train Validation Test
Provided 405 102 120 316 80 96
Humor Data GPT4 211 - - 211 - -
Riddlesense 4531 - - - - -
Table 1: Dataset Statistics, ‘-‘ means the data was not used for the stage of the task.
The task dataset and additional datasets used in our approaches are detailed in Table  1, with all datasets being in the English language. We did not perform any extra pre-processing on the original training or test data. To generate humor data, we used GPT-4 (OpenAI et al., 2024) using prompt engineering. Regarding the RiddleSense (Lin et al., 2021) dataset, which originally had five labels, we adapted it to a four-label format. This was achieved by reassigning questions with the fifth label as the correct answer to the fourth choice. Consequently, all fifth-choice answers across questions were remapped to their corresponding fourth choices, and all original fifth choices were discarded. Riddlesense and humor datasets, were selected for their similarity to the original training data, offering commonsense-defying puzzles. For details on the train-validation-test split, please refer to Table 1. We also experimented by adding SWAG (Zellers et al., 2018) and CODAH (Chen et al., 2019) datasets, but found that they reduced overall performance.
5.2 Implementation Details
The raw text was tokenized using a byte-level Byte-Pair Encoding (BPE) vocabulary with 50,257 merge rules, and inputs longer than 1024 tokens were truncated.
Our models were based on the BERT-base and DeBERTaV3 base architectures. The BERT model comprises 12 layers, 768-dimensional embeddings, and 12 attention heads, totaling 117M parameters. The DeBERTaV3 base model features 12 layers and a hidden size of 768, with 110M backbone parameters and a 128K token vocabulary introducing an additional 98M parameters in the embedding layer.
Both models were initialized with pre-trained weights in the AutoModelForMultipleChoice architecture. We conducted a random hyperparameter search, exploring batch sizes of [4, 16, 32] and learning rates of [5e-5, 1e-4, 2e-4]. The configurations yielding the highest validation accuracy were selected for each model size.
We utilized Amazon SageMaker for training, opting for the ml.p3.8xlarge instance for BERT-based approaches and the ml.p3.16xlarge instance for training our DeBERTaV3-based approaches. The training time for the BERT models with the original data was under 20 minutes, while the DeBERTa-based approaches were trained in under one hour. This efficient use of resources enabled us to achieve significant performance improvements with minimal cost and time.
6 Results
Sentence Puzzle Word Puzzle
Approaches
Original
Semantic
Context
Orig. + Sem.
Orig. + Sem. + Con.
Overall
Original
Semantic
Context
Orig. + Sem.
Orig. + Sem. + Con.
Overall
Human .907 .907 .944 .907 .889 .920 .917 .917 .917 .917 .900 .917
ChatGPT .608 .593 .679 .507 .397 .627 .561 .524 .518 .439 .292 .535
RoBERTa-L .435 .402 .464 .330 .201 .434 .195 .195 .232 .146 .061 .207
BERT-base + AMSC + train-data-wp+sp .475 .55 .5 .35 .25 .508 .281 .312 .375 .031 0 .323
BERT-base + AMMC + train-data-wp+sp .650 .625 .625 .600 .500 .600 .438 .375 .406 .344 .375 .406
DeBERTaV3 + AMMC + train-data-wp+sp .900 .900 .850 .900 .825 .883 .75 .75 .625 .719 .500 .708
DeBERTaV3 + AMMC + train-data-wp+sp + Humor + RiddleSense .925 .950 .900 .925 .875 .925 - - - - - -
DeBERTaV3 + AMMC + train-data-wp + Humor - - - - - - .844 .812 .750 .781 .594 .802
Table 2: SemEval2024 Task 9: BRAINTEASER results table, which shows the performance of different approaches on the test set. Orig. = Original, Sem. = Semantic, Con. = Context, AMSC = AutoModelForSequenceClassification, AMMC = AutoModelForMultipleChoice
In Table 2, we demonstrate the performance of our model, where the provided numbers represent the accuracy for various groups. "Original," "Semantic," and "Context" denote the original question, its semantic reconstruction, and context reconstruction, respectively. These three categories are based on instance-based accuracy, where each question is treated as a separate instance. The score reports the accuracy for both the original question and its adversarial counterparts. "Orig. + Sem." represents group-based accuracy, where the original question and its semantic reconstruction are considered and calculated together. Similarly, "Orig. + Sem. + Con." includes the previous group along with the contextual reconstruction of the original question.
In the table, "AMSC" represents AutoModelForSequenceClassification, and "AMMC" represents AutoModelForMultipleChoice. The models used are bert-base-uncased and microsoft/deberta-v3-base. The notation "train-data-wp+sp" indicates that the training data for this approach includes both sentence puzzle and word puzzle training data provided by the organizers of the task. "Humor" represents the synthetic dataset generated by prompting GPT-4, and "RiddleSense" refers to the open-source RiddleSense dataset (Lin et al., 2021). The scores of human performance and the baseline system, as provided in the original paper (Jiang et al., 2023), are depicted in gray. Scores obtained by our system are shown in black, with the best performances for each task highlighted in bold.
6.1 Subtask A : Sentence Puzzle
Initially, we trained our models only on the provided sentence puzzle dataset but soon realized that combining both the sentence puzzle and word puzzle datasets yielded better validation scores. Consequently, we used the bert-base model with AutoModelForSequenceClassification, achieving an overall accuracy of 50.8%. Given that the dataset is in a multiple-choice format, we experimented with AutoModelForMultipleChoice using the same bert model. This change significantly improved performance, increasing accuracy by 10 points to 60%. Encouraged by this, we opted for the larger DeBERTaV3 model under the AutoModelForMultipleChoice configuration. This model, combined with the original dataset, significantly boosted performance, raising overall accuracy to 88.3%. After incorporating additional datasets containing humor-style questions and the RiddleSense dataset, our best accuracy score reached 92.5%. Our approach ranked 6th among the 31 teams that participated in the task and outperformed the baseline zero shot ChatGPT by almost 50 percentage points.
6.2 Subtask B : Word Puzzle
The Word Puzzle setup followed a similar approach to the Sentence Puzzle. However, during the validation process, we found that the best-performing model was trained only on the original Word Puzzle training data combined with the humor dataset. Adding RiddleSense data and Sentence Puzzle data did not improve Word Puzzle validation scores, so we did not include those configurations in our final submission. This is a notable asymmetry: while Sentence Puzzle benefits from diverse augmentation, Word Puzzle does not. We hypothesize that this is because word-level puzzles rely on orthographic and phonetic manipulation rather than semantic reinterpretation, making sentence-oriented augmentation data less relevant. Our approach for this subtask ranked 10th among the 23 participating teams, but still outperformed the baseline zero-shot ChatGPT by approximately 40 percentage points.
6.3 Ablation Study
To isolate the contribution of each component in our system, we summarize the incremental gains on the Sentence Puzzle subtask in Table 3. Each row adds one component relative to the row above it.
The largest single gain (+28.3 points) comes from switching from BERT to DeBERTaV3, confirming the importance of the pre-trained model’s capacity. The architectural choice between sequence classification and multiple choice also yields a substantial improvement (+9.2 points), underscoring that task formulation is a meaningful design decision, not merely an implementation detail. Data augmentation provides a further 4.2-point gain on Sentence Puzzle. For Word Puzzle, the best configuration (DeBERTaV3 + AMMC + wp + Humor) achieves 80.2% overall accuracy, while adding RiddleSense or Sentence Puzzle data reduces performance.
Configuration Acc. $\Delta$
BERT + AMSC + wp + sp 50.8
BERT + AMMC wp + sp 60.0 +9.2
DeBERTaV3 + AMMC + wp + sp 88.3 +28.3
 + Humor + RS 92.5 +4.2
Table 3: Ablation on Sentence Puzzle (overall accuracy, %). RS = RiddleSense, AMSC = AutoModelForSequenceClassification, AMMC = AutoModelForMultipleChoice.
6.4 Error Analysis
We analyze the errors made by our best system to identify systematic failure patterns.
Instance-based vs. Group-based Accuracy:
A notable pattern across all configurations is the significant drop from instance-based to group-based accuracy. For example, on the Sentence Puzzle, our best system achieves 92.5% on original questions but only 87.5% on the Orig.+Sem.+Con. group metric. This indicates that while the model often answers the original question correctly, it fails on at least one adversarial variant, suggesting that its reasoning is not fully robust—it may be relying on surface patterns that break under semantic or contextual reconstruction.
Sentence Puzzle vs. Word Puzzle Gap:
Our system shows a 12.3-point gap between Sentence Puzzle (92.5%) and Word Puzzle (80.2%) overall accuracy. Sentence Puzzles require reinterpreting the semantic content of a sentence, which aligns well with the semantic representations learned by transformer models. Word Puzzles, in contrast, often require reasoning about orthographic structure (e.g., identifying that “N” is the part of “London” that is in “France”), which is not directly captured by subword tokenization. This suggests that improving Word Puzzle performance may require architectural changes or character-level augmentation strategies rather than additional semantic training data.
Effect of Augmentation on Word Puzzles:
Adding RiddleSense data, which helped the Sentence Puzzle, actually decreased Word Puzzle validation scores. RiddleSense riddles are primarily semantic in nature, which may introduce noise when the target task requires character-level or phonetic reasoning. This finding highlights that data augmentation must be carefully matched to the reasoning type required by the target task.
7 Conclusion
We presented a system for the SemEval 2024 BRAINTEASER task that achieves 92.5% accuracy on the Sentence Puzzle (6th/31 teams) and 80.2% on the Word Puzzle (10th/23 teams). Our results demonstrate three key findings: (1) the multiple-choice architecture is substantially better suited to this task than sequence classification, (2) augmenting training data with humor and riddle datasets effectively improves sentence-level lateral reasoning, and (3) data augmentation benefits do not transfer uniformly across reasoning types—word-level puzzles require different strategies. The gap between instance-based and group-based accuracy across all systems highlights that adversarial robustness remains a challenge for lateral thinking tasks. Future work should explore character-level representations or tokenization-aware approaches for word puzzles, larger-scale synthetic data generation with better quality control, and ensemble methods that combine semantic and orthographic reasoning capabilities.
8 Limitations
Our work has several limitations that should be considered when interpreting the results.
Small Training Data:
The provided training data is small (405 Sentence Puzzle and 316 Word Puzzle instances). While our augmentation strategy mitigates this to some extent, the limited data constrains the generalizability of our conclusions about which augmentation strategies are most effective.
Base-size Models Only:
We experimented exclusively with base-size models (BERT-base, DeBERTaV3-base) due to computational constraints. Larger models or recent LLM-based approaches (e.g., instruction-tuned models with in-context learning) may exhibit different augmentation dynamics and could potentially close the gap on Word Puzzles.
Synthetic Data Quality:
Our humor dataset was generated by prompting GPT-4 with relatively simple prompts and manually checked for duplicates. A more systematic generation and filtering pipeline could yield higher-quality and more diverse augmentation data. Additionally, the humor data generation process is not fully reproducible, as GPT-4 outputs are non-deterministic.
Limited Analysis of Adversarial Robustness:
While we report group-based accuracy metrics that capture robustness to semantic and context reconstruction, we do not perform a detailed analysis of which types of adversarial perturbations are most challenging. A finer-grained analysis could inform more targeted approaches to improving robustness.
Single Task Evaluation:
Our approach is evaluated solely on the BRAINTEASER task. It remains an open question whether the data augmentation strategies we propose generalize to other lateral thinking or creative reasoning benchmarks.
Appendix A Humor Dataset Details
We used the following prompts to generate the humor-style dataset. We experimented with multiple prompts and gathered all outputs in a JSON file, then analyzed them manually.
PROMPT 1 - Could you create a dataset for me that includes humor-styled questions, each with multiple choices and an answer? The dataset should be in JSON format.
PROMPT 2 - Could you create a dataset of 40 jokes for me in JSON format? Each joke should include four options and the correct answer.
PROMPT 3 - Could you generate an additional 20 jokes with multiple choices and an answer? Please ensure there are no duplicates and that none of them are the same as those previously generated.
We initially prompted GPT-4 to generate 200 questions at once, but the output contained duplicate questions after approximately 15–16 unique ones, as the model began repeating itself. We therefore used PROMPT 3 iteratively to generate data in smaller batches, checking for duplicates manually before each addition. Below are examples of the generated jokes.
"joke": "Why did the bicycle fall over?", "options": [ "A. Because it was two-tired.", "B. It had a flat.", "C. It was unbalanced.", "D. It slipped." ], "answer": "A"
"joke": "What’s orange and sounds like a parrot?", "options": [ "A. A carrot", "B. An orange bird", "C. A tangerine", "D. A flamingo" ], "answer": "A" ,